4 3 INI 1 1 1 2 3 4. 2 To find the blue shaded area above, we would calculate: Isle f(x)dx = area Where: b= f(x) = Area =
Answer #1

Soln : 3. 8-2 2 C 2 1 2 3 4 The blue region lies between x=1 and a=2. a=1 & b = 2 = f(x) Also , the blue region is under theArea () * - * ) | 8 – 4) -(4-1) ] 1) 4 - 3 4 - 13 ง
Answer #2

SOLUTION :


            b

Area = ∫  f(x) dx

           a

where, 


a = 1 , b = 2 ,


f(x) = - x / 2 + 4  (since slope is - 1/2, and y-intercept is 4)


So, 


            

Area 


   2

= ∫  (- x / 2 + 4) dx 

   1

                             2

=  [- x^2 / 4 + 4x ]

                             1


= [ (- 2^2 / 4 + 4*2) - (- 1^2 / 4 + 4*1)]


= [ 7  - 15/4 ]


= 13/4 


= 3.25   (ANSWER)



Check :


Graphically , Area = (3.5 + 3)/2 * 1 = 3.25 (same as above).

answered by: Tulsiram Garg